21x^2-16/3x^2-4 -7x=5 find or solve for the value of x in the above equation

Grade:10

1 Answers

HARISH SAHU
7 years ago
Given Equation is
21x^2 - \frac{16}{3}x^2 -4 -7x =5
both side multiply by 3
 
63x^2 -16x^2 -12-21x =15
arranging simila terms
47x^2 -21x -27 =0
after solving
x1 = 1.0135798980727626 and x2 = -0.5667713874344646.........Ans